Henry's sister is 3 years younger than Henry. The product of their ages is 180.
What are their ages?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52778)

.
15 and 12 are two factors of 180 that differ in 3 units and whose product is exactly 180.



And there are no other such positive numbers.



(But still there are negative: -15 and -12).



Notice that I presented you the MENTAL solution.



There is another way to solve the problem using equation:


x*(x+3) = 180,

x^2 + 3x - 180 = 0,

(x-12)*(x+15) = 0,


which gives the roots  x = 12 and x = -15.


Answer by MathTherapy(10551)

Henry's sister is 3 years younger than Henry. The product of their ages is 180.
What are their ages?